A circle with a diameter of 6 inches has its center at the center of a square with 6-inch sides. Find the area of the region that is inside the square and outside the square. Use 3.14 for Pi
Answer by rfer(12677)

A=3.14*3^2
A=28.26
------------
a= 6*6
a=36
-----------
36-28.26=7.74 sq in